What is the equation of this line? (-1 , 2). (4 , -1)
Answer by stanbon(75887)

What is the equation of this line? (-1 , 2). (4 , -1)
------
slope =(-1-2)/(4--1) = -3/5
--------------
Form:: y = mx + b
2 = -(3/5)*-1 + b
b = 7/5
-----
Equation:
y = (-3/5)x + (7/5)
=======================
Cheers,
Stan H.
=======================